A sly old fox ate 100 grapes in 5 days, each day eating 6 more than on the day before. How many did he eat on each of the 5 days?
Answer by checkley75(3666)

X+(X+6)+(X+12)+(X+18)+(X+24)
5X+60=100
5X=100-60
5X=40
X=40/5
X=8 ANSWER FOR THE NUMBER OF GRAPES EATEN ON THE FIRST DAY.
8+6=14 FOR THE SECOND DAY
14+6=20 FOR THE THIRD DAY
20+6=26 FOR THE FOURTH
26+6=32 FOR THE FIFTH DAY
PROOF
8+14+20+26+32=100
100=100